OpenStudy (anonymous):

i need help...it says write an equation of a parabola with the vertex at (1,0) and the directrix at x=-5....someone please help : or explain it to me

OpenStudy (anonymous):

put it in cartesian coordinates or just draw this one first... here directrix x=-5, means the parabola opens to the right....its formula is 4p(x-h)=(y-k)^2,.....vertex(1,0) ,,,h=1,k=0,,directrix=h-p,.... -5=1-p p=1+5=6 therefore 4p(x-h)=(y-k)^2 4(6)(x-1)=(y-0)^2 24(x-1)=y^2 24x-24=y^2 ans or y^2=24x-24 ans

OpenStudy (anonymous):

goodluck babygurl..